\begin{thebibliography}{99}
%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
%
%[1]
\bibitem{H1Z:2009wt}
  F.~D.~Aaron {\it et al.} [ H1 and ZEUS Collaborations ],
  %``Combined Measurement and QCD Analysis of the Inclusive e+- p Scattering Cross 
  %   Sections at HERA,''
  JHEP {\bf 1001 } (2010)  109.
  [arXiv:0911.0884 [hep-ex]].
%-----------------------------------------------------------------------------------
%
%[2]
\bibitem{PDF}
%\cite{Alekhin:2010iu}
%\bibitem{Alekhin:2010iu}
  S.~Alekhin, J.~Bl{\"u}mlein and S.~Moch,
  %``Update of the NNLO PDFs in the 3-, 4-, and 5-flavour scheme,''
  arXiv:1007.3657 [hep-ph];\\
  %%CITATION = ARXIV:1007.3657;%%
%\cite{Alekhin:2009ni}
%\bibitem{Alekhin:2009ni}
  S.~Alekhin, J.~Bl\"umlein, S.~Klein and S.~Moch,
  %``The 3-, 4-, and 5-flavor NNLO Parton from Deep-Inelastic-Scattering Data
  %and at Hadron Colliders,''
  Phys.\ Rev.\  D {\bf 81} (2010) 014032
  [arXiv:0908.2766 [hep-ph]];\\
  %%CITATION = PHRVA,D81,014032;%%
%\cite{Blumlein:2006be}
%\bibitem{Blumlein:2006be}
  J.~Bl{\"u}mlein, H.~B{\"o}ttcher and A.~Guffanti,
  %``Non-singlet QCD analysis of deep inelastic world data at O(alpha(s)**3),''
  Nucl.\ Phys.\  B {\bf 774} (2007) 182
  [arXiv:hep-ph/0607200];
  %%CITATION = NUPHA,B774,182;%%
%\cite{Blumlein:2004ip}
%\bibitem{Blumlein:2004ip}
%  J.~Blumlein, H.~Bottcher and A.~Guffanti,
  %``Non-singlet QCD analysis of the structure function F2 in 3-loops,''
  Nucl.\ Phys.\ Proc.\ Suppl.\  {\bf 135} (2004) 152
  [arXiv:hep-ph/0407089];\\
  %%CITATION = NUPHZ,135,152;%%
%\cite{Gluck:2006yz}
%\bibitem{Gluck:2006yz}
  M.~Gl{\"u}ck, E.~Reya and C.~Schuck,
  %``Non-singlet QCD analysis of F2(x,Q**2) up to NNLO,''
  Nucl.\ Phys.\  B {\bf 754} (2006) 178
  [arXiv:hep-ph/0604116];\\
  %%CITATION = NUPHA,B754,178;%%
%\cite{JimenezDelgado:2008hf}
%\bibitem{JimenezDelgado:2008hf}
  P.~Jimenez-Delgado and E.~Reya,
  %``Dynamical NNLO parton distributions,''
  Phys.\ Rev.\  D {\bf 79} (2009) 074023
  [arXiv:0810.4274 [hep-ph]];\\
  %%CITATION = PHRVA,D79,074023;%%
%\cite{Martin:2009iq}
%\bibitem{Martin:2009iq}
  A.~D.~Martin, W.~J.~Stirling, R.~S.~Thorne and G.~Watt,
  %``Parton distributions for the LHC,''
  Eur.\ Phys.\ J.\  C {\bf 63} (2009) 189
  [arXiv:0901.0002 [hep-ph]].
  %%CITATION = EPHJA,C63,189;%%
%----------------------------------------------------------------------------------- 
% 
%[3] 
%\cite{Buza:1995ie}
\bibitem{Buza:1995ie}
  M.~Buza, Y.~Matiounine, J.~Smith, R.~Migneron and W.~L.~van Neerven,
  %``Heavy quark coefficient functions at asymptotic values $Q~2 \gg m~2$,''
  Nucl.\ Phys.\  B {\bf 472} (1996) 611.
  %[hep-ph/9601302].
  %%CITATION = NUPHA,B472,611;%%
%-----------------------------------------------------------------------------------
%
%[4]
%\cite{Vermaseren:2005qc}
\bibitem{Vermaseren:2005qc}
  J.~A.~M.~Vermaseren, A.~Vogt and S.~Moch,
  %``The third-order QCD corrections to deep-inelastic scattering by photon
  %exchange,''
  Nucl.\ Phys.\  B {\bf 724} (2005) 3.
  %[hep-ph/0504242].
  %%CITATION = NUPHA,B724,3;%%
%-----------------------------------------------------------------------------------
%
%[5]
%\cite{Bierenbaum:2009mv}
\bibitem{Bierenbaum:2009mv}
  I.~Bierenbaum, J.~Bl{\"u}mlein and S.~Klein,
  %{\sf Mellin Moments of the {$O(\alpha_s^3$)} Heavy Flavor Contributions to
  %unpolarized Deep-Inelastic Scattering at $Q^2 \gg m^2$ and Anomalous
  %Dimensions}, 
  Nucl.\ Phys.\  B {\bf 820} (2009) 417.
  % [hep-ph/0904.3563].
  %%CITATION = ARXIV:0904.3563;%%
%-----------------------------------------------------------------------------------
%
%[6]
\bibitem{Steinhauser:2000ry}
  M.~Steinhauser,
  %``MATAD: A program package for the computation of massive tadpoles,''
  Comput.\ Phys.\ Commun.\  {\bf 134} (2001) 335
  [arXiv:hep-ph/0009029].
  %%CITATION = CPHCB,134,335;%%
%-----------------------------------------------------------------------------------
%
%[7]
\bibitem{TOL}
%\cite{Buza:1996wv}
%\bibitem{Buza:1996wv}
  M.~Buza, Y.~Matiounine, J.~Smith and W.~L.~van Neerven,
  %``Charm electroproduction viewed in the variable-flavour number scheme
  %versus fixed-order perturbation theory,''
  Eur.\ Phys.\ J.\  C {\bf 1} (1998) 301,
  [hep-ph/9612398];\\
  %%CITATION = EPHJA,C1,301;%%
%\cite{Buza:1996xr}
%\bibitem{Buza:1996xr}
  M.~Buza, Y.~Matiounine, J.~Smith and W.~L.~van Neerven,
  %``O (alpha-s**2) corrections to polarized heavy flavor production at Q**2 is >> m**2,''
  Nucl.\ Phys.\  {\bf B485 } (1997)  420,
  [hep-ph/9608342];\\
%\cite{Bierenbaum:2007qe}
%\bibitem{Bierenbaum:2007qe}
  I.~Bierenbaum, J.~Bl\"umlein and S.~Klein,
  %``Two-loop massive operator matrix elements and unpolarized heavy flavor
  %production at asymptotic values Q**2 >> m**2,''
  Nucl.\ Phys.\  B {\bf 780} (2007) 40;
  [hep-ph/0703285];
  %%CITATION = NUPHA,B780,40;%%
%\cite{Bierenbaum:2007dm}
%\bibitem{Bierenbaum:2007dm}
%  I.~Bierenbaum, J.~Bl\"umlein and S.~Klein,
  %``Calculation of massive 2-loop operator matrix elements with outer gluon
  %lines,''
  Phys.\ Lett.\  B {\bf 648} (2007) 195,
   [hep-ph/0702265];
  %%CITATION = PHLTA,B648,195;%%
%
%\cite{Bierenbaum:2009zt}
%\bibitem{Bierenbaum:2009zt}
%  I.~Bierenbaum, J.~Bl{\"u}mlein and S.~Klein,
  %``The Gluonic Operator Matrix Elements at O(\alpha_s^2) for DIS Heavy Flavor
  %Production,''
  Phys.\ Lett.\  B {\bf 672} (2009) 401,
  [hep-ph/0901.0669];
  %%CITATION = PHLTA,B672,401;%% 
%\cite{Bierenbaum:2007pn}
%\bibitem{Bierenbaum:2007pn}
%  I.~Bierenbaum, J.~Bl\"umlein, S.~Klein,
  %``Two-loop massive operator matrix elements for polarized and unpolarized deep-inelastic scattering,''
  [arXiv:0706.2738 [hep-ph]]
               and in preparation;\\
% 
%\cite{Bierenbaum:2008yu} 
%\bibitem{Bierenbaum:2008yu}
  I.~Bierenbaum, J.~Bl\"umlein, S.~Klein and C.~Schneider,
  %``Two--Loop Massive Operator Matrix Elements for Unpolarized Heavy Flavor
  %Production to $O(\epsilon),''
  Nucl.\ Phys.\  B {\bf 803} (2008) 1,
  [hep-ph/0803.0273].
  %%CITATION = NUPHA,B803,1;%%
%-----------------------------------------------------------------------------------
%
%[8]
%\cite{Blumlein:2006mh}
\bibitem{Blumlein:2006mh}
  J.~Bl\"umlein, A.~De Freitas, W.~L.~van Neerven and S.~Klein,
  %``The longitudinal heavy quark structure function F(L)(Q anti-Q) in the
  %region Q**2 >> m**2 at O(alpha(s)**3),''
  Nucl.\ Phys.\  B {\bf 755} (2006) 272.
  % [hep-ph/0608024].
  %%CITATION = NUPHA,B755,272;%%
%-----------------------------------------------------------------------------------
%
%[9]
%\cite{Blumlein:2009rg
\bibitem{Blumlein:2009rg}
  J.~Bl\"umlein, S.~Klein and B.~T\"odtli,
  %``$O(\alpha_s^2)$ and $O(\alpha_s^3)$ Heavy Flavor Contributions to
  %Transversity at $Q^2 \gg m^2$,''
  Phys.\ Rev.\  {\bf D 80} (2009) 094010.
  %[arXiv:0909.1547 [hep-ph]].
  %%CITATION = PHRVA,D80,094010;%%
%-----------------------------------------------------------------------------------
%
%[10]
\bibitem{Ablinger1}
J.~Ablinger, J.~Bl{\"u}mlein,  S.~Klein, C.~Schneider and F.~Wi\ss{}brock in preparation;\\
%-----------------------------------------------------------------------------------
%
%[11]
\bibitem{Ablinger:2010ha}
  J.~Ablinger, I.~Bierenbaum, J.~Bl{\"u}mlein, A.~Hasselhuhhn, S.~Klein, C.~Schneider and F.~Wi\ss{}brock,
  %``Heavy Flavor DIS Wilson coefficients in the asymptotic regime,''
  arXiv:1007.0375 [hep-ph].
%-----------------------------------------------------------------------------------
%
%[12]
\bibitem{SIGMA}
C. Schneider, J. Symbolic Comput. 43 (2008) 611,
\newblock [arXiv:0808.2543v1]; Ann. Comb. 9 (2005) 75;
J. Differ. Equations Appl. 11 (2005) 799; Ann. Comb.  (2009) to appear,
[arXiv:0808.2596]; {Proceedings of the Conference on Motives, Quantum Field Theory, and
Pseudodifferential Operators}, To appear in the Mathematics Clay Proceedings,
2010; S\'em.~Lothar. Combin. 56 (2007) 1, Article B56b,  Habilitationsschrift JKU Linz (2007)
and references therein;\\
%\cite{Ablinger:2010pb}
%\bibitem{Ablinger:2010pb}
  J.~Ablinger, J.~Bl{\"u}mlein, S.~Klein and C.~Schneider,
  %``Modern Summation Methods and the Computation of 2- and 3-loop Feynman
  %Diagrams,''
  arXiv:1006.4797 [math-ph].
  %%CITATION = ARXIV:1006.4797;%%
%-----------------------------------------------------------------------------------
%
%[13]
\bibitem{HSUM}
%\cite{Blumlein:1998if}
%\bibitem{Blumlein:1998if}
  J.~Bl{\"u}mlein and S.~Kurth,
  %``Harmonic sums and Mellin transforms up to two-loop order,''
  Phys.\ Rev.\  D {\bf 60} (1999) 014018
  [arXiv:hep-ph/9810241];\\
  %%CITATION = PHRVA,D60,014018;%%
%\cite{Vermaseren:1998uu}
%\bibitem{Vermaseren:1998uu}
  J.~A.~M.~Vermaseren,
  %``Harmonic sums, Mellin transforms and integrals,''
  Int.\ J.\ Mod.\ Phys.\  A {\bf 14} (1999) 2037
  [arXiv:hep-ph/9806280].
  %%CITATION = IMPAE,A14,2037;%%
%-----------------------------------------------------------------------------------
%
%[14]
\bibitem{ANDIM1}
%\cite{Larin:1993vu}
%\bibitem{Larin:1993vu}
  S.~A.~Larin, T.~van Ritbergen and J.~A.~M.~Vermaseren,
  %``The Next Next-To-Leading QCD Approximation For Nonsinglet Moments Of Deep
  %Inelastic Structure Functions,''
  Nucl.\ Phys.\  B {\bf 427} (1994) 41;\\
  %%CITATION = NUPHA,B427,41;%%
%\cite{Larin:1996wd}
%\bibitem{Larin:1996wd}
  S.~A.~Larin, P.~Nogueira, T.~van Ritbergen and J.~A.~M.~Vermaseren,
  %``The 3-loop QCD calculation of the moments of deep inelastic  structure
  %functions,''
  Nucl.\ Phys.\  B {\bf 492} (1997) 338;
  [hep-ph/9605317];\\
  %%CITATION = NUPHA,B492,338;%%
%\cite{Retey:2000nq}
%\bibitem{Retey:2000nq}
  A.~Retey and J.~A.~M.~Vermaseren,
  %``Some higher moments of deep inelastic structure functions at
  %next-to-next-to leading order of perturbative QCD,''
  Nucl.\ Phys.\  B {\bf 604} (2001) 281;
  [hep-ph/0007294];\\
  %%CITATION = NUPHA,B604,281;%%
%\cite{Blumlein:2004xt}
%\bibitem{Blumlein:2004xt}
  J.~Bl\"umlein and J.~A.~M.~Vermaseren,
  %``The 16th moment of the non-singlet structure functions F2(x,Q**2) and
  %F(L)(x,Q**2) to O(alpha(s)**3),''
  Phys.\ Lett.\  B {\bf 606} (2005) 130;
  [hep-ph/0411111];\\
  %%CITATION = PHLTA,B606,130;%%
\bibitem{Gracey:2003yrxGracey:2003mrxGracey:2006zrxGracey:2006ah}
  J. A. Gracey,
     Nucl. Phys. {\bf B662}, 247 (2003)~;
     Nucl. Phys. {\bf B667}, 242 (2003)~;
     JHEP {\bf 10}, 040 (2006)~;
     Phys. Lett. {\bf B643}, 374 (2006)~;
%-----------------------------------------------------------------------------------
%
%[15]
\bibitem{ANDIM2}
%\bibitem{Gracey:1993nn}
  J. A. Gracey,
  Phys. Lett. {\bf B 322} (1994) 141;\\ 
%\cite{Moch:2004pa}
%\bibitem{Moch:2004pa}
  S.~Moch, J.~A.~M.~Vermaseren and A.~Vogt,
  %``The three-loop splitting functions in QCD: The non-ninglet case,''
  Nucl.\ Phys.\  B {\bf 688} (2004) 101;
  [hep-ph/0403192].\\
  %%CITATION = NUPHA,B688,101;%%
%\cite{Vogt:2004mw}
%\bibitem{Vogt:2004mw}
  %A.~Vogt, S.~Moch and J.~A.~M.~Vermaseren,
  %``The three-loop splitting functions in QCD: The singlet case,''
  Nucl.\ Phys.\  B {\bf 691} (2004) 129;
  [hep-ph/0404111].
  %%CITATION = NUPHA,B691,129;%%
%-----------------------------------------------------------------------------------
%
%[16]
\bibitem{BBK1}
I. Bierenbaum, J. Bl{\"u}mlein, and S. Klein, in preparation.
%-----------------------------------------------------------------------------------
\end{thebibliography}

\documentclass{PoS}

\title{\vspace*{-2mm} {\tiny DESY 10--122 \hfill SFB/CPP-10-74 \hfill TTK-10-40 \hfill IFIC-10-23}
\\
\boldmath Logarithmic $O(\alpha_s^3)$ contributions to the DIS Heavy Flavor Wilson 
Coefficients 
at $Q^2 \gg m^2$}

\ShortTitle{Logarithmic $O(\alpha_s^3)$ contributions}

\author{Isabella Bierenbaum\thanks{Supported in part by Generalitat Valenciana  Grant No. 
PROMETEO/2008/69.}\\
        Instituto de F\'{i}sica Corpuscular, CSIC-Universitat de Val\`{e}ncia, Apartado de Correos 
22085, 
E--46071 Valencia, Spain\\
        E-mail: \email{Isabella.Bierenbaum@ific.uv.es}}

\author{\speaker{Johannes Bl\"umlein}\thanks{Supported in part by SFB-TR/9 and EU TMR network 
HEPTOOLS.}\\
        Deutsches Elektronen--Synchrotron, DESY, Platanenallee 6, D--15638 Zeuthen, Germany\\
        E-mail: \email{Johannes.Bluemlein@desy.de}}

\author{Sebastian Klein\thanks{Supported in part by SFB-TR/9.}\\
        Institut f\"ur theoretische Teilchenphysik und Kosmologie, RWTH Aachen University, D--52056 
Aachen, Germany\\
        E-mail: \email{sklein@physik.rwth-aachen.de}}

\abstract{The logarithmic contributions to the massive twist-2 operator matrix elements for 
deep-inelastic scattering are calculated to $O(\alpha_s^3)$for general values of the Mellin 
variable $N$.}

\FullConference{XVIII International Workshop on Deep-Inelastic Scattering and Related Subjects, DIS 2010\\
		April 19-23, 2010\\
		Firenze, Italy}

\newcommand{\N}{\nonumber}

\newcommand{\Ctil}{\tilde{C}}

\begin{document}
\section{Introduction}

\noindent
The heavy flavor contributions to deep-inelastic scattering (DIS) are rather large in the small $x$ 
region.
With the current DIS data a precision of $\sim 1\%$ is reached~\cite{H1Z:2009wt} for $F_2(x,Q^2)$. This 
requires to describe the heavy flavor corrections to 3-loop order, to perform a consistent next-to-next-to 
leading order (NNLO) analysis, to measure the strong coupling constant $\alpha_s(M_Z^2)$ and the 
unpolarized twist-2 parton distribution functions at highest precision possible,~cf.~\cite{PDF}.
In the region $Q^2/m^2 \geq 10$ one may compute all contributions except 
the power suppressed terms, $\propto (m^2/Q^2)^k, k\geq 1$ using the factorization 
theorem given in Ref.~\cite{Buza:1995ie}. Here, the massive Wilson coefficients factorize into massive
operator matrix elements (OMEs), $A_{ij}$, and the massless Wilson coefficients, which are known to 
$O(\alpha_s^3)$~\cite{Vermaseren:2005qc}. The $O(\alpha_s^3)$ Mellin moments of the massive operator matrix 
elements up to $N = 10 ... 14$, depending on the process, were computed in \cite{Bierenbaum:2009mv}.
The calculation was performed relating the moments of the massive operator matrix elements    
to massive tadpoles and using {\tt MATAD}~\cite{Steinhauser:2000ry}.
In Ref.~\cite{Bierenbaum:2009mv} also the complete renormalization for a single massive quark has been 
derived. Different 
other contributions, needed in the renormalization process, were computed at general values of $N$ in 
Refs.~\cite{TOL}. For the structure 
function $F_L(x,Q^2)$ the asymptotic corrections to $O(\alpha_s^3)$ are known for general 
values of $N$~\cite{Blumlein:2006mh}. They are, however, only applicable at scales $Q^2/m^2 \geq 800$.
For transversity the matrix elements were computed for general $N$ at 2-loop order and a series of
moments at 3-loop order in \cite{Blumlein:2009rg}. 
Very recently, the general $N$ results at 
$O(\alpha_s^3)$ for $F_2(x,Q^2)$ for the color coefficients $\propto n_f$ have been 
completed~\cite{Ablinger1,Ablinger:2010ha}. These computations use modern summation technologies encoded
in the package {\tt Sigma}~\cite{SIGMA} and the results can be expressed in terms of nested harmonic sums
\cite{HSUM}. 
From the single pole terms in the massive computations of
Refs.~\cite{Bierenbaum:2009mv,Blumlein:2009rg,Ablinger1,Ablinger:2010ha}
the corresponding contributions to the 3--loop anomalous dimensions were derived, either for the respective 
moments \cite{ANDIM1}, or at general values of $N$~\cite{ANDIM2}.

In this note we report on the logarithmic $O(\alpha_s^3)$ contributions to the massive operator matrix 
elements,~cf. also \cite{BBK1}. They are known for general values of $N$ and depend on the 3-loop anomalous 
dimensions and massive OMEs up to $O(\alpha_s^2)$. 
\section{The heavy flavor Wilson coefficients in the asymtotic region}

\noindent
The heavy flavor correction to the structure function $F_2(x,Q^2)$ with $n_f$ massless and one
heavy flavor reads,~\cite{Bierenbaum:2009mv}~:
\begin{eqnarray}
      \label{eqF2}
       F_{(2,L)}^{Q\overline{Q}}(x,n_f,Q^2,m^2) &=&
       \sum_{k=1}^{n_f}e_k^2\Biggl\{
                   L_{q,(2,L)}^{\sf NS}\left(x,n_f+1,\frac{Q^2}{m^2}
                                                ,\frac{m^2}{\mu^2}\right)
                \otimes
                   \Bigl[f_k(x,\mu^2,n_f)+f_{\overline{k}}(x,\mu^2,n_f)\Bigr]
\N\\ &&\hspace{-35mm}
               +\frac{1}{n_f}\Bigl[L_{q,(2,L)}^{\sf PS}\left(x,n_f+1,\frac{Q^2}{m^2}
                                                ,\frac{m^2}{\mu^2}\right)
                \otimes
                   \Sigma(x,\mu^2,n_f)
               + L_{g,(2,L)}^{\sf S}\left(x,n_f+1,\frac{Q^2}{m^2}
                                                 ,\frac{m^2}{\mu^2}\right)
                \otimes
                   G(x,\mu^2,n_f)
                             \Bigr] \Biggr\}
\N\\ &&\hspace{-35mm}
+ e_Q^2\Biggl[
                   H_{q,(2,L)}^{\sf PS}\left(x,n_f+1,\frac{Q^2}{m^2}
                                        ,\frac{m^2}{\mu^2}\right)
                \otimes
                   \Sigma(x,\mu^2,n_f)
                  +H_{g,(2,L)}^{\sf S}\left(x,n_f+1,\frac{Q^2}{m^2}
                                           ,\frac{m^2}{\mu^2}\right)
                \otimes
                   G(x,\mu^2,n_f)
                                  \Biggr]~,\N\\
\end{eqnarray}
with boundaries for the Mellin integral $[x(1+4m^2/Q^2), 1]$, and $e_i$ the quark charges.
Here the different Wilson coefficients are denoted by $L_i, H_i$ in case the photon couples to
a light (L) or the heavy (H) quark. For $Q^2 \gg m^2$ they can be expressed in terms of the
massive OMEs $A_{ij}$ and the massless Wilson coefficients $C_j$. To $O(a_s^3)$ they read ($a_s = 
\alpha_s/(4\pi))$
\begin{eqnarray}
\label{eqWIL1}
     L_{q,(2,L)}^{\sf NS}(n_f+1) &=&
     a_s^2 \Bigl[A_{qq,Q}^{(2), {\sf NS}}(n_f+1) \delta_2 +
     \hat{C}^{(2), {\sf NS}}_{q,(2,L)}(n_f)\Bigr]
     \N\\
     &+&
     a_s^3 \Bigl[A_{qq,Q}^{(3), {\sf NS}}(n_f+1) \delta_2
     +  A_{qq,Q}^{(2), {\sf NS}}(n_f+1) C_{q,(2,L)}^{(1), {\sf NS}}(n_f+1)
     + \hat{C}^{(3), {\sf NS}}_{q,(2,L)}(n_f)\Bigr]  
      \label{eqWIL2}
\nonumber\\
      L_{q,(2,L)}^{\sf PS}(n_f+1) &=&
     a_s^3 \Bigl[~A_{qq,Q}^{(3), {\sf PS}}(n_f+1)~\delta_2
     +  A_{gq,Q}^{(2)}(n_f)~~n_f\Ctil_{g,(2,L)}^{(1)}(n_f+1) 
     + n_f \hat{\Ctil}^{(3), {\sf PS}}_{q,(2,L)}(n_f)\Bigr]
         \label{eqWIL3}
\nonumber\\
      L_{g,(2,L)}^{\sf S}(n_f+1) &=&
     a_s^2 A_{gg,Q}^{(1)}(n_f+1)n_f \Ctil_{g,(2,L)}^{(1)}(n_f+1)
+
      a_s^3 \Bigl[~A_{qg,Q}^{(3)}(n_f+1)~\delta_2 \N\\ &&
     +  A_{gg,Q}^{(1)}(n_f+1)~~n_f\Ctil_{g,(2,L)}^{(2)}(n_f+1)
     +  A_{gg,Q}^{(2)}(n_f+1)~~n_f\Ctil_{g,(2,L)}^{(1)}(n_f+1)
     \N\\ && 
     +  ~A^{(1)}_{Qg}(n_f+1)~~n_f\Ctil_{q,(2,L)}^{(2), {\sf PS}}(n_f+1)
     + n_f \hat{\Ctil}^{(3)}_{g,(2,L)}(n_f)\Bigr]~,
\nonumber\\
     H_{q,(2,L)}^{\sf PS}(n_f+1)
     &=& a_s^2 \Bigl[~A_{Qq}^{(2), {\sf PS}}(n_f+1)~\delta_2
     +~\Ctil_{q,(2,L)}^{(2), {\sf PS}}(n_f+1)\Bigr]
+ a_s^3 \Bigl[~A_{Qq}^{(3), {\sf PS}}(n_f+1)~\delta_2 
\N\\ &&
     +~\Ctil_{q,(2,L)}^{(3), {\sf PS}}(n_f+1) 
     + A_{gq,Q}^{(2)}(n_f+1)~\Ctil_{g,(2,L)}^{(1)}(n_f+1)
     \N\\ &&
+ A_{Qq}^{(2), {\sf PS}}(n_f+1)~C_{q,(2,L)}^{(1), {\sf NS}}(n_f+1)
        \Bigr]~,       \label{eqWIL4}
\nonumber\\
     H_{g,(2,L)}^{\sf S}(n_f+1) &=& a_s \Bigl[~A_{Qg}^{(1)}(n_f+1)~\delta_2
     +~\Ctil^{(1)}_{g,(2,L)}(n_f+1) \Bigr]   + a_s^2 \Bigl[~A_{Qg}^{(2)}(n_f+1)~\delta_2
\N\\ &&
     +~A_{Qg}^{(1)}(n_f+1)~C^{(1), {\sf NS}}_{q,(2,L)}(n_f+1)
     +~A_{gg,Q}^{(1)}(n_f+1)~\Ctil^{(1)}_{g,(2,L)}(n_f+1)
     \N\\ &&
     +~\Ctil^{(2)}_{g,(2,L)}(n_f+1) \Bigr]
     +~a_s^3 \Bigl[~A_{Qg}^{(3)}(n_f+1)~\delta_2
     +~A_{Qg}^{(2)}(n_f+1)~C^{(1), {\sf NS}}_{q,(2,L)}(n_f+1)
     \N\\ &&
     +~A_{gg,Q}^{(2)}(n_f+1)~\Ctil^{(1)}_{g,(2,L)}(n_f+1)
     +~A_{Qg}^{(1)}(n_f+1)\Bigl\{
     C^{(2), {\sf NS}}_{q,(2,L)}(n_f+1)
     \N\\ && 
     +~\Ctil^{(2), {\sf PS}}_{q,(2,L)}(n_f+1)\Bigr\}
     +~A_{gg,Q}^{(1)}(n_f+1)~\Ctil^{(2)}_{g,(2,L)}(n_f+1)
     +~\Ctil^{(3)}_{g,(2,L)}(n_f+1) \Bigr]~, 
\N\\
\label{eqWIL5}
\end{eqnarray}
with $\delta_2 = 0 (1)$ for $F_L (F_2)$ and $\hat{f}(n_f) = f(n_f+1) - f(n_f), \tilde{f}(n_f) = 
f(n_f)/n_f$.
The massive OMEs depend on the ratio $m^2/\mu^2$, while the scale ratio of the massless 
Wilson coefficients is $\mu^2/Q^2$. The latter are pure functions of the momentum fraction $z$, or the
Mellin variable $N$, if one sets $\mu^2 = Q^2$. The massive OMEs obey then the general structure
\begin{eqnarray}
\label{eq:log}
A^{(3)}_{ij}\left(\frac{m^2}{Q^2}\right)  = 
  a^{(3),3}_{ij} \ln^3\left(\frac{m^2}{Q^2}\right)
+ a^{(3),2}_{ij} \ln^2\left(\frac{m^2}{Q^2}\right)
+ a^{(3),1}_{ij} \ln\left(\frac{m^2}{Q^2}\right)
+ a^{(3),0}_{ij}~.
\end{eqnarray}
\section{The matrix element $A_{Qg}^{(3)}(N)$}

\noindent
In the following we present, as an example, the logarithmic expansion coefficients of Eq.~(\ref{eq:log}), 
$a^{(3),k}, k \geq 1$, for the massive OME $A_{Qg}^{(3)}(N)$ in the $\overline{\rm MS}$--scheme. They are 
given by~:
%
%
{\small
 \begin{eqnarray}
   a_{Qg}^{(3),3}&=&\frac{8(N^2+N+2)T_F}{9N(N+1)(N+2)}\Biggl[
     T_Fn_f\Biggl(
          C_F\Biggl(
                \frac{P_1}{(N-1)N^2(N+1)^2(N+2)}
               -4S_{1}
             \Biggr)
 \N\\ && \hspace{-10mm}
         +C_A\Biggl(
                4S_{1}
               -\frac{8(N^2+N+1)}{(N-1)N(N+1)(N+2)}
             \Biggr)
           \Biggr)
    -8T_F^2
    +C_A^2 \Biggl(
             -\frac{(11N^4+22N^3-59N^2-70N-48)S_{1}}{(N-1)N(N+1)(N+2)}
 \N\\ && \hspace{-10mm}
             -12S_{1}^2
             +\frac{2(N^2+N+1)(11N^4+22N^3-35N^2-46N-24)}{(N-1)^2N^2(N+1)^2(N+2)^2}
           \Biggr)
    +C_AT_F\Biggl(
             -\frac{56(N^2+N+1)}{(N-1)N(N+1)(N+2)}
 \N\\ && \hspace{-10mm}
             +28S_{1}
           \Biggr)
    +C_F^2\Biggl(
             -3\frac{(3N^2+3N+2)^2}{4N^2(N+1)^2}
             +\frac{6S_{1}(3N^2+3N+2)}{N(N+1)}
             -12S_{1}^2
          \Biggr)
    +C_FT_F\Biggl(
             -16S_{1}
 \N\\ && \hspace{-10mm}
             +\frac{2P_2}{(N-1)N^2(N+1)^2(N+2)}
           \Biggr)
    +C_AC_F\Biggl(
              24S_{1}^2
             -\frac{(N^2+N+6)(7N^2+7N+4)S_{1}}{(N-1)N(N+1)(N+2)}
 \N\\ && \hspace{-10mm}
             -\frac{(3N^2+3N+2)(11N^4+22N^3-59N^2-70N-48)}{4(N-1)N^2(N+1)^2(N+2)}
          \Biggr)
                                                    \Biggr]~.\label{AQg33}
 \end{eqnarray}
 }
%
%
{\small
 \begin{eqnarray}
  a_{Qg}^{(3),2}&=&4T_F^2n_f\Biggl(
    C_F\Biggl(
         -\frac{4(N^2+N+2)}{3N(N+1)(N+2)}\Bigl(S_{2}+S_{1}^2\Bigr)
         +\frac{8(5N^3+8N^2+19N+6)S_{1}}{9N^2(N+1)(N+2)}
 \N\\ && \hspace{-10mm}
         -\frac{P_3}{9(N-1)N^4(N+1)^4(N+2)^3}
       \Biggr)
   +C_A\Biggl(
         -\frac{8(5N^4+20N^3+47N^2+58N+20)S_{1}}{9N(N+1)^2(N+2)^2}
 \N\\ && \hspace{-10mm}
         +\frac{4(N^2+N+2)}{3N(N+1)(N+2)}\Bigl(2S_{-2}+S_{2}+S_{1}^2\Bigr)
         -\frac{2P_4}{9(N-1)N^2(N+1)^3(N+2)^3}
       \Biggr)
                         \Biggr)
  +2C_A^2T_F\Biggl(
 \N\\ && \hspace{-10mm}
               \frac{8(N^2+N+2)}{N(N+1)(N+2)}
                \Bigl(2S_{-2,1}-S_{1}^3-3S_{2}S_{1}-S_{-3}-4S_{-2}S_{1}-S_{3}\Bigr)
              -\frac{2P_5S_{1}^2}{3(N-1)N^2(N+1)^2(N+2)^2}
 \N\\ && \hspace{-10mm}
              +\frac{4P_6S_{1}}{9(N-1)^2N^3(N+1)^3(N+2)^3}
              -\frac{4(N^2+N+2)(11N^4+22N^3-59N^2-70N-48)S_{-2}}{3(N-1)N^2(N+1)^2(N+2)^2}
 \N\\ && \hspace{-10mm}
              +\frac{8P_7}{9(N-1)^2N^4(N+1)^4(N+2)^4}
              -\frac{2(N^2+N+2)(11N^4+22N^3-83N^2-94N-72)S_{2}}{3(N-1)N^2(N+1)^2(N+2)^2}
            \Biggr)
 \N\\ && \hspace{-10mm}
  +4C_AT_F^2\Biggl(
               \frac{4(N^2+N+2)}{N(N+1)(N+2)}\Bigl(S_{1}^2+S_{2}+2S_{-2}\Bigr)
              +\frac{8(5N^4+20N^3-N^2-14N+20)S_{1}}{9N(N+1)^2(N+2)^2}
 \N\\ && \hspace{-10mm}
              -\frac{2P_8}{9(N-1)N^3(N+1)^3(N+2)^3}
           \Biggr)
  +2C_F^2T_F\Biggl(
               \frac{8(N^2+N+2)}{N(N+1)(N+2)}
                \Bigl(3S_{2}S_{1}-S_{1}^3+4S_{-2}S_{1}+2S_{3}-4S_{-2,1}\Bigr)
 \N\\ && \hspace{-10mm}
              -\frac{16(N^2+N+2)S_{-2}}{N^2(N+1)^2(N+2)}
              -\frac{6(N^2+N+2)(3N^2+3N+2)S_{2}}{N^2(N+1)^2(N+2)}
              +\frac{2(3N^4+14N^3+43N^2+48N+20)S_{1}^2}{N^2(N+1)^2(N+2)}
 \N\\ && \hspace{-10mm}
              -\frac{4P_9S_{1}}{N^3(N+1)^3(N+2)}
              +\frac{P_{10}}{2N^4(N+1)^4(N+2)}
              +\frac{16(N^2+N+2)S_{-3}}{N(N+1)(N+2)}
            \Biggr)
  +4C_FT_F^2\Biggl(
 \N\\ && \hspace{-10mm}
               \frac{4(N^2+N+2)}{3N(N+1)(N+2)}\Bigl(S_{2}-3S_{1}^2\Bigr)
              +\frac{8(5N^3+14N^2+37N+18)S_{1}}{9N^2(N+1)(N+2)}
              -\frac{P_{11}}{9(N-1)N^4(N+1)^4(N+2)^3}
            \Biggr)
 \N\\ && \hspace{-10mm}
  +2C_FC_AT_F\Biggl(
                \frac{4(N^2+N+2)}{N(N+1)(N+2)}
                 \Bigl(4S_{1}^3-2S_{3}+4S_{-2,1}-2S_{-3}-3S_{-2}\Bigr)
 \N\\ && \hspace{-10mm}
               +\frac{4P_{12}S_{1}^2}{3(N-1)N^2(N+1)^2(N+2)^2}
               -\frac{4P_{13}S_{1}}{9(N-1)N^3(N+1)^3(N+2)^3}
 \N\\ && \hspace{-10mm}
               +\frac{P_{14}}{18(N-1)N^3(N+1)^3(N+2)^3}
               +\frac{4(N^2+N+2)(N^4+2N^3+8N^2+7N+18)S_{2}}{3(N-1)N^2(N+1)^2(N+2)^2}
            \Biggr)~.\label{AQg32}
 \end{eqnarray}
 }
%
%
{\small
 \begin{eqnarray}
  a_{Qg}^{(3),1}&=&
    \frac{1}{2}\hat{\gamma}_{qg}^{(2)}(n_f)
   -\frac{n_f}{2}\hat{\tilde{\gamma}}_{qg}^{(2)}(n_f)
   +4T_F^2n_f\Biggl(
       C_F\Biggl(
             \frac{4(N^2+N+2)}{9N(N+1)(N+2)}
              \Bigl(4S_{3}-S_{1}^3-3S_{2}S_{1}\Bigr)
            +\frac{4(3N+2)S_{1}^2}{3N^2(N+2)}
 \N\\ && \hspace{-10mm}
            +\frac{4(N^4-N^3-20N^2-10N-4)S_{1}}{3N^2(N+1)^2(N+2)}
            +\frac{2P_{15}}{3(N-1)N^5(N+1)^5(N+2)^4}
            +\frac{4P_{16}S_{2}}{3(N-1)N^3(N+1)^3(N+2)^2}
          \Biggr)
 \N\\ && \hspace{-10mm}
      +C_A\Biggl(
             \frac{4(N^2+N+2)}{9N(N+1)(N+2)}
              \Bigl( S_{1}^3+9S_{2}S_{1}+6S_{-3}+12S_{-2}S_{1}+8S_{3}
                    -12S_{-2,1}\Bigr)
            -\frac{4P_{17}S_{1}}{3N(N+1)^3(N+2)^3}
 \N\\ && \hspace{-10mm}
            -\frac{4(N^3+8N^2+11N+2)S_{1}^2}{3N(N+1)^2(N+2)^2}
            +\frac{4P_{18}}{3(N-1)N^4(N+1)^4(N+2)^4}
            -\frac{4P_{19}S_{2}}{3(N-1)N^2(N+1)^2(N+2)^2}
 \N\\ && \hspace{-10mm}
            +\frac{16(N^2-N-4)S_{-2}}{3(N+1)^2(N+2)^2}
          \Biggr)
           \Biggr)
   +2C_A^2T_F\Biggl(
                \frac{8(N^2+N+2)}{3N(N+1)(N+2)}
                  \Bigl( 12S_{-2,1}S_{1}-S_{1}^4-9S_{2}S_{1}^2-8S_{3}S_{1}
                        -6S_{-3}S_{1}
 \N\\ && \hspace{-10mm}
                        -12S_{-2}S_{1}^2\Bigr)
               -\frac{2P_{20}S_{1}^3}{9(N-1)N^2(N+1)^2(N+2)^2}
               +\frac{2P_{21}S_{1}^2}{3(N-1)N^2(N+1)^3(N+2)^3}
 \N\\ && \hspace{-10mm}
               -\frac{2P_{22}S_{1}}{3(N-1)N^4(N+1)^4(N+2)^4}
               -\frac{2P_{23}S_{2}S_{1}}{(N-1)N^2(N+1)^2(N+2)^2}
               -\frac{2P_{24}}{3(N-1)^2N^5(N+1)^5(N+2)^5}
 \N\\ && \hspace{-10mm}
               +\frac{4(N^2+N+2)(11N^4+22N^3-35N^2-46N-24)}{9(N-1)N^2(N+1)^2(N+2)^2}\Bigl(6S_{-2,1}-4S_{3}-3S_{-3}\Bigr)
 \N\\ && \hspace{-10mm}
               -\frac{8(N^2-N-4)(11N^4+22N^3-35N^2-46N-24)S_{-2}}{3(N-1)N(N+1)^3(N+2)^3}
               -\frac{8P_{25}S_{-2}S_{1}}{3(N-1)N^2(N+1)(N+2)^2}
 \N\\ && \hspace{-10mm}
               +\frac{2(11N^4+22N^3-35N^2-46N-24)P_{19}S_{2}}{3(N-1)^2N^3(N+1)^3(N+2)^3}
             \Biggr)
   +4C_AT_F^2\Biggl(
                \frac{8(N^2+N+2)}{9N(N+1)(N+2)}
                  \Bigl( S_{1}^3+8S_{3}-12S_{-2,1}
 \N\\ && \hspace{-10mm}
                        +9S_{2}S_{1}+12S_{-2}S_{1}+6S_{-3}\Bigr)
               -\frac{8(N^3+8N^2+11N+2)S_{1}^2}{3N(N+1)^2(N+2)^2}
               -\frac{8P_{19}S_{2}}{3(N-1)N^2(N+1)^2(N+2)^2}
 \N\\ && \hspace{-10mm}
               +\frac{2P_{26}}{27(N-1)N^4(N+1)^4(N+2)^4}
               +\frac{32(N^2-N-4)S_{-2}}{3(N+1)^2(N+2)^2}
               +\frac{4P_{27}S_{1}}{27N(N+1)^3(N+2)^3}
             \Biggr)
   +2C_F^2T_F\Biggl(
 \N\\ && \hspace{-10mm}
                \frac{8(N^2+N+2)}{3N(N+1)(N+2)}
                 \Bigl(4S_{3}S_{1}-S_{1}^4-3S_{2}S_{1}^2\Bigr)
               +\frac{2(3N^4+42N^3+107N^2+92N+28)S_{1}^3}{3N^2(N+1)^2(N+2)}
 \N\\ && \hspace{-10mm}
               +\frac{2P_{28}S_{1}^2}{N^3(N+1)^2(N+2)}
               +\frac{2P_{29}S_{1}}{N^4(N+1)^4(N+2)}
               +\frac{2(7N^4+74N^3+79N^2-12N-4)S_{2}S_{1}}{N^2(N+1)^2(N+2)}
 \N\\ && \hspace{-10mm}
               -\frac{8(N^2+N+2)(3N^2+3N+2)S_{3}}{3N^2(N+1)^2(N+2)}
               -\frac{2(3N^2+3N+2)(N^4+17N^3+17N^2-5N-2)S_{2}}{N^3(N+1)^3(N+2)}
 \N\\ && \hspace{-10mm}
               -\frac{P_{30}}{N^5(N+1)^5(N+2)}
            \Biggr)
   +4C_FT_F^2\Biggl(
                \frac{8(N^2+N+2)}{9N(N+1)(N+2)}
                 \Bigl(4S_{3}-S_{1}^3-3S_{2}S_{1}\Bigr)
 \N\\ && \hspace{-10mm}
               +\frac{P_{31}}{3(N-1)N^5(N+1)^5(N+2)^2}
               +\frac{8(N^4-N^3-20N^2-10N-4)S_{1}}{3N^2(N+1)^2(N+2)}
               +\frac{8(N^4+17N^3+17N^2-5N-2)S_{2}}{3N^2(N+1)^2(N+2)}
\N\\ && \hspace{-10mm}
                +\frac{8(3N+2)S_{1}^2}{3N^2(N+2)}
             \Biggl)
   +2C_FC_AT_F\Biggl(
                 \frac{8(N^2+N+2)}{3N(N+1)(N+2)}
                  \Bigl( 2S_{1}^4+12S_{2}S_{1}^2+4S_{3}S_{1}-12S_{-2,1}S_{1}
                        +6S_{-3}S_{1}
\N\\ && \hspace{-10mm}
                        +12S_{-2}S_{1}^2\Bigr)
                +\frac{4P_{32}S_{1}^3}{9(N-1)N^2(N+1)^2(N+2)^2}
                -\frac{4P_{33}S_{1}^2}{3(N-1)N^3(N+1)^3(N+2)^3}
\N\\ && \hspace{-10mm}
                -\frac{4P_{34}S_{1}}{3(N-1)N^4(N+1)^4(N+2)^4}
                -\frac{4P_{35}S_{2}S_{1}}{3(N-1)N^2(N+1)^2(N+2)^2}
                -\frac{P_{36}}{3(N-1)N^4(N+1)^5(N+2)^4}
\N\\ && \hspace{-10mm}
                -\frac{8P_{37}S_{-2}S_{1}}{N^2(N+1)^2(N+2)^2}
                +\frac{4(N^2+N+2)(3N^2+3N+2)}{N^2(N+1)^2(N+2)}
                  \Bigl(2S_{-2,1}-S_{-3}\Bigr)
\N\\ && \hspace{-10mm}
                -\frac{8(N^2-N-4)(3N^2+3N+2)S_{-2}}{N(N+1)^3(N+2)^2}
                -\frac{2P_{38}S_{2}}{3(N-1)N^3(N+1)^3(N+2)^2}
\N\\ && \hspace{-10mm}
                -\frac{8(N^2+N+2)(29N^4+58N^3-41N^2-70N-48)S_{3}}{9(N-1)N^2(N+1)^2(N+2)^2}
            \Biggr)~.\label{AQg31}
 \end{eqnarray}
 }

\normalsize
\noindent
Here, $S_{\vec{a}} \equiv S_{\vec{a}}(N)$, $P_k$ denote some polynomials in $N$, 
cf.~\cite{BBK1}, and $\gamma_{ij}^{(2)}$ are the 3-loop anomalous dimensions. 
The expansion coefficients given above depend on harmonic sums up weight {\sf w = 3}. 
Numerical studies show, that within the kinematic region of HERA the constant
terms to (\ref{eq:log}) are as important as the logarithmic contributions. 
Further details will be given in \cite{BBK1}.

\end{document}